Germaine Semmes plans to buy a home in 17 years and will need another $120,240.00 before xhe can do so. Given an interest rate of 3.00% compounded monthly, how much money does xhe need to set aside today? a. $84,333.92 b. $72,249.56 c. $119,815.43 d. $72,747.18
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

120,240=P*(1+0.03/12)^(12*17)

P=120,240/(1+0.03/12)^(12*17)

=120,240*0.600877879

=$72,249.56(Approx).